# Unveiling the Potential of iMarkers: Invisible Fiducial Markers for Advanced Robotics

### v2

@@ 6.1 Comparative Insights
Considering the sensor-related factors shown in Fig. 9-b, the performance scores can be studied in five aspects. In “design complexity,” the static single-vision setup is the most straightforward, requiring only a static polarizer attached to the lens. The dynamic single-vision setup has a moderate complexity by incorporating a function generator to control polarization. Dual-vision configurations represent the most complex designs, with the cube beamsplitter offering more reliable light division than the plate one, which requires precise alignment for optimal beam splitting. The design complexity directly influences the “sensor cost.” The static single-vision setup offers the most economical solution, requiring only a polarizer, while the dynamic single-vision configuration incurs additional costs due to its polarization control components. Dual-vision setups are generally more expensive. They require two cameras, a beamsplitter (either cube or plate), and a structural mounting cage to ensure precise alignment and stability.
@@
Regarding the “field of view,” single-vision setups provide a broader theoretical FoV (e.g., $87^{\circ}\times 58^{\circ}$ in the RealSense™ D435, shown in Fig. 6-c). However, in dual-vision setups, the beamsplitter’s physical dimensions and the supporting cage partially obstruct the cameras’ FoV, reducing the effective viewing area. For instance, the configuration shown in Fig. 6-c achieves an effective FoV of $57^{\circ}\times 54.5^{\circ}$. Another essential factor is “weight”; the static single-vision setup is the lightest, whereas the dynamic single-vision setup adds approximately $30~{}gr$ due to the function generator and dynamic polarizer. Thus, single-vision setups are more feasible for minimizing payloads, such as in drones, as dual-vision configurations are heavier. The combined weight of dual-vision setup components typically ranges from $150$ to $350~{}gr$, with the plate beamsplitter potentially reducing the overall sensor weight. Finally, considering the noise levels in the processed output, dual-vision setups produce clearer results, isolating only the CSR regions of the iMarkers. In contrast, the dynamic single-vision setup may introduce additional noise due to motion and illumination fluctuations during “temporal” subtraction. The static single-vision setup retains the highest noise pixels (i.e., pixels that do not contribute to iMarker detection), including the filtered areas with the detected iMarkers.
@@ 6.2 Applicability
Considering the unique characteristics of iMarkers, they serve as potential alternatives in scenarios reliant on printed markers. By safely replacing printed markers, iMarkers can preserve unobtrusiveness and aesthetic integrity while delivering full functionality for robotics and AR systems. Their obscure nature makes them appropriate for surroundings with unwanted visual clutter, such as museums, hospitals, offices, etc. Fig. 10 depicts the real-world benchmarking of iMarkers in Visual Simultaneous Localization and Mapping (VSLAM), showcasing the use of a visible-range iMarker to facilitate indoor map reconstruction enriched with semantic information [20]. iMarkers can also be suitable for autonomous robots functioning in retail stores and smart warehousing, where numerous markers are needed to augment objects. Another potential use case includes object grasping in cluttered environments, where iMarkers can supply accurate localization cues to enhance robotic manipulation accuracy. iMarkers can also be strategically employed to seamlessly augment difficult-to-detect objects, like transparent surfaces (e.g., windows and glass block walls), where the transparency poses considerable challenges to computer vision and robot perception systems. Likewise, reflective surfaces, such as mirrors or chrome/metallic objects, often cause distortions or reflections, confusing perception algorithms and negatively impacting object recognition or localization accuracy. In these scenarios, iMarkers can augment such objects by providing reliable information while preserving visual appeal. This allows robots and AR devices to effectively see and interact with visually challenging entities, ensuring reliable operation and versatility.
@@
6.3 Limitations
Given the ongoing nature of this research and the absence of a rapid iMarker fabrication process so far, the experiments remain constrained under varying illumination, weather conditions (for outdoor applications), and environmental factors (e.g., dirt). Additionally, the reflectivity of UV-range CSRs needs to be enhanced to ensure that all UV-range iMarkers within the FoV are reliably detected when illuminated by a broad-field light source.
@@ Acknowledgments
The authors would like to thank Marco Giberna for his insightful discussions and contributions to the experiments conducted in this work.

### v3

@@ 6.1 Comparative Insights
Considering the sensor-related factors shown in Fig. 10-b, the performance scores can be studied in five aspects. In “design complexity,” the static single-vision setup is the most straightforward, requiring only a static polarizer attached to the lens. The dynamic single-vision setup has moderate complexity by incorporating a function generator to control polarization. Dual-vision configurations represent the most complex designs, with the cube beamsplitter offering more reliable light division than the plate one, which requires precise alignment for optimal beam splitting. The design complexity directly influences the “sensor cost.” The static single-vision setup offers the most economical solution, requiring only a polarizer, while the dynamic single-vision configuration incurs additional costs due to its polarization control components. Dual-vision setups are generally more expensive. They require two cameras, a beamsplitter (either cube or plate), and a structural mounting cage to ensure precise alignment and stability.
@@
Regarding the “field of view,” single-vision setups provide a broader theoretical FoV (e.g., $87^{\circ}\times 58^{\circ}$ in the RealSense™ D435, shown in Fig. 6-c). However, in dual-vision setups, the beamsplitter’s physical dimensions and the supporting cage partially obstruct the cameras’ FoV, reducing the effective viewing area. For instance, the configuration shown in Fig. 6-c achieves an effective FoV of $57^{\circ}\times 54.5^{\circ}$. Another essential factor is “weight”; the static single-vision setup is the lightest, whereas the dynamic single-vision setup adds approximately 30gr30~gr due to the function generator and dynamic polarizer. Thus, single-vision setups are more feasible for minimizing payloads, such as in drones, as dual-vision configurations are heavier. The combined weight of dual-vision setup components typically ranges from $150$ to 350gr350~gr, with the plate beamsplitter potentially reducing the overall sensor weight. Finally, considering the noise levels in the processed output, dual-vision setups produce clearer results, isolating only the CSR regions of the iMarkers. In contrast, the dynamic single-vision setup may introduce additional noise due to motion and illumination fluctuations during “temporal” subtraction. The static single-vision setup retains the highest noise pixels (i.e., pixels that do not contribute to iMarker detection), including the filtered areas with the detected iMarkers.
When choosing among various sensor configurations, a selection parameter is the trade-off between “detection robustness” and “clutter rejection,” as each configuration’s detection pipeline can be tailored to optimize performance for specific application scenarios. Dual-vision setups rely on spatial frame subtraction followed by thresholding, producing cleaner outputs by isolating CSR-coated iMarker regions while effectively suppressing non-marker backgrounds thanks to the differential visibility of CSRs between the two sensor views. However, they are more sensitive to artifacts caused by sensor noise, motion, illumination variations, and frame misalignment. In contrast, single-vision setups may retain more background clutter (noise pixels) alongside detected iMarkers. Accordingly, enhancing robustness and output quality are tackled differently considering the employed sensor variant. While strategies such as adaptive thresholding, active frame registration, and noise modeling can better aid dual-vision setup algorithms, single-vision systems can benefit from lightweight morphological filtering to enhance detection quality without compromising system simplicity.
@@ 6.2 Applicability
Considering the unique characteristics of iMarkers, they serve as potential alternatives in scenarios reliant on printed markers. In this regard, the obscure nature of iMarkers makes them appropriate for surroundings with unwanted visual clutter, such as museums, hospitals, offices, etc. By safely replacing printed markers, iMarkers can preserve unobtrusiveness and aesthetic integrity while delivering full functionality for robotics and AR systems. It should be noted that various AR applications, particularly those deployable on handheld mobile devices, rely exclusively on RGB cameras and frequently utilize fiducial landmarks to support interactive functionalities like user engagement. This aligns well with the iMarker detection procedure proposed in this paper; specifically the single-vision setup which requires minimal hardware modifications to enable compatibility with standard AR-capable vision sensors.
Additionally, as the iMarker detection pipeline relies on simple computer vision techniques (e.g., thresholding and frame subtraction), it introduces minimal computational overhead and operates efficiently even on a CPU-based system, without requiring hardware acceleration. Thus, the core innovation of the proposed system lies in tightly coupling hardware and software, where the sensor hardware captures frames with filtered views on CSR-coated regions of iMarkers and the software extracts and reveals these regions before passing the result to standard fiducial marker libraries (e.g., ArUco or AprilTag).
Fig. 11 depicts the real-world benchmarking of iMarkers in Visual Simultaneous Localization and Mapping (VSLAM), showcasing the use of a visible-range iMarker to facilitate indoor map reconstruction enriched with semantic information [20]. iMarkers can also be suitable for autonomous robots functioning in retail stores and smart warehousing, where numerous markers are needed to augment objects. Another potential use case includes object grasping in cluttered environments, where iMarkers can supply accurate localization cues to enhance robotic manipulation accuracy. iMarkers can also be strategically employed to seamlessly augment difficult-to-detect objects, like transparent surfaces (e.g., windows and glass block walls), where the transparency poses considerable challenges to computer vision and robot perception systems. Likewise, reflective surfaces, such as mirrors or chrome/metallic objects, often cause distortions or reflections, confusing perception algorithms and negatively impacting object recognition or localization accuracy. In these scenarios, iMarkers can augment such objects by providing reliable information while preserving visual appeal. This allows robots and AR devices to effectively see and interact with visually challenging entities, ensuring reliable operation and versatility.
@@
6.3 Limitations and Fabrication Considerations
Given the ongoing nature of this research and the absence of a rapid iMarker fabrication process so far, the experiments remain constrained under varying illumination, weather conditions (for outdoor applications), and environmental factors (e.g., dirt). In terms of environmental robustness, iMarkers are expected to exhibit similar vulnerability to dirt and occlusion as conventional printed markers, since both rely on visual input for marker detection and pose estimation. In other words, the marker recognition pipeline (ArUco in this work) processes either standard RGB frames of normal cameras or the binary outputs of the iMarker detector algorithm, without introducing additional sensitivity. Regarding sunlight and outdoor use, UV-range iMarkers may experience reduced contrast under ambient sunlight, potentially leading to false positives due to uncontrolled UV reflections. Accordingly, the reflectivity of UV-range CSRs needs to be enhanced to ensure that all UV-range iMarkers are reliably detected when illuminated by the light source. Similarly, IR-range iMarkers could be influenced by environmental IR sources. For these cases in real-world deployments, additional filtering strategies or controlled illumination may be required to ensure reliable iMarker detection.
The scalability of iMarker fabrication for mass deployment is still an active investigation topic. Currently, production involves several manual steps, including microfluidic droplet generation, annealing, polymerization, and sequential washing, which reduce throughput and introduce variability in marker quality. To overcome this, we are developing an integrated, automated production platform that combines all stages into a continuous process with real-time computer monitoring. This system is expected to streamline manufacturing, improve consistency, and enable large-scale high-throughput fabrication suitable for real-world robotics applications.
@@ Acknowledgments
The authors would like to thank Marco Giberna and Tadej Emersic for their insightful assistance during experiments conducted in this work.